OpenStudy (anonymous):

What is the value of x in the equation −x = 3 − 4x + 6?

OpenStudy (anonymous):

A. -9 B. -3 C. 3 D. 9

Mehek (mehek14):

first add 4x to -x 4x+(-x)=?

OpenStudy (anonymous):

-4x?

Mehek (mehek14):

no add 4 + (-1)

OpenStudy (anonymous):

-5

Mehek (mehek14):

that's the same as 4-1

OpenStudy (anonymous):

ok

Mehek (mehek14):

so what do you get 4-1

OpenStudy (anonymous):

idk i suck at math

Mehek (mehek14):

you don't know 4 - 1???

OpenStudy (anonymous):

oh 3 hahha

Mehek (mehek14):

ok so you have 3x = 3 + 6

Mehek (mehek14):

now add 3 + 6

OpenStudy (anonymous):

yep, so 9

Mehek (mehek14):

ok so \(3x=9\) to get x by itself, divide \(\dfrac{9}{3}\)

OpenStudy (anonymous):

its 3

Mehek (mehek14):

correct

OpenStudy (anonymous):

so the answers 3! thanks for your help!